The width of a pool table is x feet, and the length is 4 feet longer than the width. Find polynomials A(x) and P(x) that represent the area and perimeter of the pool table. Find A(4) and P(4).
Answer by ankor@dixie-net.com(15667)

The width of a pool table is x feet, and the length is 4 feet longer than the width. Find polynomials A(x) and P(x) that represent the area and perimeter of the pool table.
:
We have:
x = the width
(x+4) = length
:
Area
A(x) = x(x+4)
A(x) = x^2 + 4x
:
Perimeter
P(x) = 2(x+4) + 2x
P(x) = 2x + 8 + 2x
P(x) = 4x + 8
:
:
Find A(4)
A(4) = 4^2 + 4(4)
A(4) = 16 + 16
A(4) = 32
:
and P(4).
P(4) = 4(4) + 8
P(4) = 16 + 8
P(4) = 24